OpenStudy (anonymous):

show the series with term ln r/r^2 converges

OpenStudy (anonymous):

i.e. \[\ln r \over r^2\]

OpenStudy (anonymous):

i figure if this term is \[u _{n}\] then if i can find \[a_{n}\] that is < than \[u _{n}\] that also converges i will be able to show it; however 1/r^2 doesn't get me there and I can't show it for 1/r^3, hence stumped

OpenStudy (anonymous):

i've seen a few questions with ln r in numerator, so i'm guessing this is a common thing

OpenStudy (anonymous):

\[\lim_{n \rightarrow \infty} \frac{\ln}{n ^{2}}\]\[\lim_{n \rightarrow \infty} \frac{1/n}{2n}\] \[\lim_{n \rightarrow \infty} \frac{-1/n ^{2}}{2}\] =0

OpenStudy (anonymous):

what is ln/n^2

OpenStudy (anonymous):

I made a mistake, it is lnn/n^2

OpenStudy (anonymous):

i don't follow your logic there, are you comparing with the term 1/2n^2

OpenStudy (anonymous):

do you know l`hopital`s rule

OpenStudy (anonymous):

no

OpenStudy (anonymous):

its not mentioned in my further math book

OpenStudy (anonymous):

i can only assume its possible without it

OpenStudy (anonymous):

ok it is says that if there is \[\frac{\infty}{\infty}\] you can derivative proportion and denominator

OpenStudy (anonymous):

oh,thats interesting

OpenStudy (anonymous):

i'll google it

OpenStudy (anonymous):

thanks for your help, odd that the book would ask the qu and no mention of the rule

OpenStudy (anonymous):

i think the answer is to compare with r^0.5

OpenStudy (anonymous):

i.e. as r^0.5/r^2 converges ( this is the same as 1/r^1.5) and ln r < r^0.5 for all r>0